OpenStudy (anonymous):

Find the volume of the region between the graph of f(x,y)=25−x2−y2 and the xyplane

OpenStudy (zarkon):

I would switch to polar coordinates

OpenStudy (zarkon):

\[\int\limits_{-5}^{5}\int\limits_{-\sqrt{25-y^2}}^{\sqrt{25-y^2}}(25-x^2-y^2)dxdy\] \[=\int\limits_{0}^{2\pi}\int\limits_{0}^{5}(25-r^2)r\,drd\theta\] \[=\frac{625\pi}{2}\]